40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3242     Entry time: Sun Jul 18 20:50:03 2010
Author: rana 
Type: Configuration 
Category: IOO 
Subject: MC TRANS optics changed 

To make the beam on the MC trans camera bigger, I removed the lens + ND filter that was in that path.

The camera was getting the transmission through a BS1-33 (33% reflector). The reflection went to the TRANS QPD. I changed

the R=33% into an HR mirror (Y1-45P) so now the camera has a nice beam. The QPD was now saturating so I put a ND06 into that path

so now the TRANS_SUM is ~4.5-5 V when the MC is aligned.

The MC was also misaligned and failing to lock all weekend (why??) so I aligned the MC mirrors to get it to acquire again. Since we want to

collect MC seismic data, please make sure the MC is locked and running after finishing with your various MC or PSL work (this means YOU).

ELOG V3.1.3-